The initial rate data for the reaction 2N2O5(g) → 4NO2(g) + O2(g) is shown in the following table.

            Experiment    [N2O5](M)        Rate (M/s)

                1         1.28 × 102         22.5

                2         2.56 x 102         45.0

a. Determine the order of reaction of the given reactant.

b. Determine the value of the rate constant for this reaction.


1
Expert's answer
2022-03-17T15:01:27-0400

(a) rate law expression can be given as follows


"rate=k[N_2O_5]^a ......................(i)"


"22.5=k[1.28\times 10^2]^a ......................(i)"


"45=k[2.56\times 10^2]^a ......................(ii)"


dividing (ii) by (i)


"2=[2]^a"


On compairing both

value of a is 1

order of reaction

for finding rate constant use equation 1

(b)

"45=k[2.56\times 10^2]^1"


"k= 17.578 \times 10^{-2} s^{-1}"


"k= 1.76 \times 10^{-1} s^{-1}"